OpenStudy (lisa123):

factor 2(x + 1)(x − 3)2 − 3(x + 1)2(x − 3)

OpenStudy (anonymous):

did u mean 2(x+1)(x-3)^2 - 3(x+1)^2(x-3)

OpenStudy (lisa123):

yes sorry

OpenStudy (anonymous):

(x-3)(-x^2+2x-9)

OpenStudy (lisa123):

how did you get that

OpenStudy (anonymous):

We need to factor the GCF (Greatest Common Factor). The resulting term is a product of the GCF and the original expression divided by the GCF.

OpenStudy (anonymous):

The GCF is (x+1)(x-3)

OpenStudy (lisa123):

okay then im left with 2(x-3) -3(x+1)?

OpenStudy (lisa123):

@Giantclan

OpenStudy (campbell_st):

so that becomes 2x - 3x -6 - 3 = -x - 9 so if you take the GCF (x +1)(x -3) have (x +1)(x-3)(-x - 9) = -(x + 1)(x-3)(x + 9)

OpenStudy (lisa123):

thank you!